The intensity, at the
central maxima (0) in a
Young's double slit set up
is 10. If the distance OP
x above P) equals one
third of the fringe width
of the pattern, show that
the intensity, at point P
would be equal to I0/4.
Distance of the screen
from the slits is D.
Distance between the
two slits is d and opex

Grade:12

1 Answers

Arun
4 years ago
The intensity at a distance y from central maxima is given by,
I = I0cos2(πdy/(λD))
Now y = x = Fringe width/3 = λD/3d
Hence,
I = I0cos2(πdy/(λD)) = I0cos2(π/3) = 
 = I0/4